MARC (marc_d):

In a certain country 1/2 of 5 = 3. If the same proportion holds, what is the value of 1/3 of 10 ?

MARC (marc_d):

@imqwerty just a riddle xD

TheSmartOne (thesmartone):

uhh mhmm (5 * 2) - 2 = 3 so (10 * 3) - 3 = 27 so would 1/3 of 10 b 27? .-.

MARC (marc_d):

nope..good try.. xD

zepdrix (zepdrix):

4? Hmm this is strange :d

MARC (marc_d):

Lol,u got the answer.. btw,can u show ur working pls ? ^_^

TheSmartOne (thesmartone):

or is it (1/2 * 5) + 1/2 = 2.5 + 0.5 = 3 (1/3 * 10) + 1/3 = 10/3 + 1/3 = 11/3 which is 3.6667

zepdrix (zepdrix):

\[\large\rm \frac{\frac12of~5}{\frac13of~10}=\frac{3}{x}\]And then just cross multiply. I feel like I shouldn't be allowed to do that though in this weird country >.< Because if 1/2 of 5 is 3, then the normal rules of multiplication shouldn't hold. I guess I'm overthinking it lolol

MARC (marc_d):

nice try @TheSmartOne but the answer is 4.. let's hear how @zepdrix got that answer..

OpenStudy (a_burning_masquerade):

2

zepdrix (zepdrix):

\[\large\rm \frac{5}{2}x=\frac{10}{3}\cdot3\]And then I did some division multiplication stuffs... :p

MARC (marc_d):

nice.. :P here is another working..

TheSmartOne (thesmartone):

that looked so confusing (and I'm not in the mood to understand what it's trying to do) that I like zeppy's way better xD

zepdrix (zepdrix):

Hmm neato :D

TheSmartOne (thesmartone):

I don't understand this step mhmm http://prntscr.com/c2aquw

MARC (marc_d):

that's what i'm thinking too @TheSmartOne.. :P didn't know there is another solution for that problem..Thanks to @zepdrix maths problem--> a lot of solution can be used to solve that particular question.. Thank you @zepdrix @TheSmartOne ^_^

OpenStudy (faiqraees):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @TheSmartOne I don't understand this step mhmm http://prntscr.com/c2aquw \(\color{#0cbb34}{\text{End of Quote}}\) What's wrong there?

TheSmartOne (thesmartone):

ohhh, they simplified the fraction from 2/6 Nevermind, I got it xD

MARC (marc_d):

xD

OpenStudy (welshfella):

1/2 * 5 = 3 5 = 6 10 = 12 1/3 * 12 = 4

OpenStudy (welshfella):

so 1/3 of 10 = 4

TheSmartOne (thesmartone):

But their answer seems flawed, imo sometimes they do this  (like they did in the first second step) and then they do this in the last 2 steps